# AlphaViT: A Flexible Game-Playing AI for Multiple Games and Variable Board Sizes

### v2

@@ 5.2 Variation of Elo rating over training iteration
Figure 4 shows the Elo ratings of the AlphaViT variants (AlphaViT, AlphaViD, AlphaVDA) and AlphaZero over multiple iterations for large and small board configurations in three games: Connect4, Gomoku, and Othello. The left column shows the results for the large board configurations (standard sizes for Connect4, Gomoku, and Othello). Here, the Elo ratings for single-game-trained agents (AlphaViT L4 LB, AlphaViD L1 LB, AlphaVDA L1 LB), multi-game-trained agents (AlphaViT L4 Multi, AlphaViD L1 Multi, AlphaVDA L1 Multi), and AlphaZero were evaluated at iterations 1, 2, 4, 6, 8, 10, 20, 40, 60, 80, 100, 200, 300, 400, 500, 600, 700, 800, 900, 1000, 1100, 1200, 1300, 1400, 1500, 1600, 1700, 1800, 1900, 2000, 2100, 2200, 2300, 2400, 2500, 2600, 2700, 2800, 2900, and 3000. The right column displays the performance on small board configurations (Connect4 5x4, Gomoku 6x6, and Othello 6x6), with evaluations of single-game-trained models (AlphaViT L4 SB, AlphaViD L1 SB, AlphaVDA L1 SB) and AlphaZero at iterations 1, 2, 4, 6, 8, 10, 20, 40, 60, 80, 100, 200, 300, 400, 500, 600, 700, 800, 900, and 1000.
For games with large boards, all single-game-trained models (AlphaViT L4 LB, AlphaViD L1 LB, AlphaVDA L1 LB) show rapid improvements in Elo ratings, reaching high levels around 300 iterations for Othello and 500 iterations for Gomoku. For Connect4, the Elo ratings continue to improve steadily, with gains even after 1000 iterations. The multi-game-trained models (AlphaViT L4 Multi, AlphaViD L1 Multi, and AlphaVDA L1 Multi) exhibit a similar trend in their Elo ratings to the single-game-trained models. Their performance across the different games shows the flexibility of these agents. After the initial rapid improvement phase, the agents show only minor gains, and their Elo ratings fluctuate.
For small boards, AlphaViT L4 SB, AlphaViD L1 SB, and AlphaVDA L1 SB show very rapid improvements in Elo ratings. After peaking, they reach stable Elo ratings for Connect4 5x4 and Gomoku 6x6. However, for Othello 6x6, the Elo ratings exhibit more fluctuations even after reaching high-performance levels. Because the Elo ratings of the agents rapidly increase and become stable, the evaluation iterations are set to be fewer than those used for larger boards.
These results show that AlphaViT, AlphaViD, and AlphaVDA perform well on small boards, achieving high and stable Elo ratings. However, for games with large boards, agents’ Elo ratings exhibit larger fluctuations. This suggests that further optimizations, such as increasing the depth of the DNNs or adjusting hyperparameters, may be necessary to enhance game-playing skills on larger boards.
@@
Figure 4: Elo rating progression over training iterations for AlphaViT, AlphaViD, and AlphaVDA in large- and small-board games. The left column shows results for large board configurations (Connect4, Gomoku, and Othello), while the right column displays results for their small board configurations (Connect4 5x4, Gomoku 6x6, and Othello 6x6). Solid lines represent single-game-trained agents, whereas dashed lines indicate multi-game-trained agents.
@@ 5.3 Effect of transformer encoder depth on performance
In this subsection, we evaluate the performance of agents with deep encoders: AlphaViT L8, AlphaViD L5, and AlphaVDA L5. Figure 5 shows the variations in Elo ratings for these agents over the iterations. The opponents and evaluation conditions are consistent with those used in previous experiments. For the large board configurations (Connect4, Gomoku, and Othello), the Elo ratings stabilize at about 1000-2000 iterations. In contrast, the Elo ratings stabilize at around 100 iterations for Connect4 5x4 and 200 iterations for Gomoku 6x6. For Othello 6x6, the Elo ratings quickly reach a high level but exhibit some fluctuations.
Table 4 provides the mean Elo ratings for each agent, calculated from 2100 to 3000 iterations for the large board configurations. Figure 6 displays the ratio of the mean Elo rating of the agents with deep encoders to that with shallow encoders. The results demonstrate that deeper architectures generally enhance the performance across all games. Notably, AlphaViD Multi and AlphaVDA Multi show significant improvements, particularly for Gomoku, which reach ratios of 1.220 and 1.165, respectively. These agents also improve the performance for Connect4, with ratios of 1.130 and 1.143, and for Othello, with ratios of 1.105 and 1.146. The single-game-trained AlphaViD and AlphaVDA exhibit significant gains for Connect4 and Othello, while their improvements for Gomoku are more modest. In contrast, the single- and multi-game-trained AlphaViTs show weaker improvements than the other agents. These results suggest that increasing the depth of transformer encoder layers enhances the agent’s performance across different board configurations and game types, with a particularly positive impact on AlphaViD and AlphaVDA.
Figure 5: Elo rating progression over training iterations for AlphaViT, AlphaViD, and AlphaVDA with deep encoder in large- and small-board games. The left column shows results for large board configurations (Connect4, Gomoku, and Othello), while the right column displays results for their small board configurations (Connect4 5x4, Gomoku 6x6, and Othello 6x6). Solid lines represent single-game-trained agents, whereas dashed lines indicate multi-game-trained agents.
@@
Game Connect4 Gomoku Othello
AlphaViT L4 LB 1966 1753 2007
AlphaViT L8 LB 2028 1862 2046
AlphaViD L1 LB 1814 1656 1756
AlphaViD L5 LB 1902 1704 1928
AlphaVDM L1 LB 1802 1570 1807
AlphaVDM L5 LB 1996 1602 1992
AlphaViT L4 Multi 1897 1917 1964
AlphaViT L8 Multi 2015 2037 1965
AlphaViD L1 Multi 1690 1611 1711
AlphaViD L5 Multi 1909 1966 1891
AlphaVDA L1 Multi 1721 1583 1686
AlphaVDA L5 Multi 1968 1844 1933
AlphaZero 2068 2187 1890
Figure 6: The ratio of Elo ratings between deep and shallow DNNs for AlphaViT, AlphaViD, and AlphaVDM across Connect4, Gomoku, and Othello. Ratios greater than 1.0 indicate superior performance of agents with deep DNNs compared to those with shallow DNNs.
@@ 5.4 Effect of fine-tuning from small board games
Figure 7 illustrates the Elo ratings of AlphaViT L8, AlphaViD L5, and AlphaVDA L5 with fine-tuned DNNs and those without fine-tuning (initialized with random weights) across three games: Connect4, Gomoku, and Othello. The fine-tuned DNNs are initialized using weights from the single-game-trained DNNs trained on the small-board configuration at 200 iterations. For example, the fine-tuned DNN for Connect4 is initialized with weights from the DNN trained on Connect4 5x4. In contrast, agents with non-fine-tuned DNNs are initialized with random weights.
@@
The results demonstrate that the agents with the fine-tuned DNNs consistently achieve higher Elo ratings than those with randomly initialized DNNs, particularly for Gomoku and Othello. The performance difference is pronounced in the early training iterations, with fine-tuned DNNs exhibiting rapid improvements and stabilization for Gomoku and Othello. For Gomoku, the advantage of fine-tuning is obvious, as the agents with the fine-tuned DNNs consistently outperform those with the non-fine-tuned DNNs throughout the training process. Some signs of overfitting appear in AlphaViD L5 and AlphaVDA L5 for Gomoku, as their Elo ratings gradually decline from approximately 1000 to 2000 iterations. However, their performance improves after 2000 iterations. In Connect4, the impact of fine-tuning is less significant, with no substantial difference in Elo ratings between agents using fine-tuned and non-fine-tuned DNNs.
@@
Figure 7: Elo rating progression over training iterations of the agents (AlphaViT L8, AlphaViD L5, and AlphaVDA L5) with finetuned and randomly initialized DNNs for three board games (Connect4, Gomoku, and Othello). The three columns show the results for AlphaViT L8, AlphaViD L5, and AlphaVDA L5, respectively. The three rows show the results for Connect4, Gomoku, and Othello.
6 Conclusion and discussion
@@
In this paper, we present AlphaViT, AlphaViD, and AlphaVDA, which are novel game-playing AI agents designed to overcome the limitations of AlphaZero using Vision Transformers (ViT). Unlike AlphaZero, which is restricted to fixed board sizes, our proposed agents demonstrate adaptability, effectively handle variations in board size, and exhibit flexibility across games. Furthermore, these agents can simultaneously train on and play multiple games, such as Connect4, Gomoku, and Othello, within a single shared neural network. The performance of these multitask agents surpasses that of traditional game AI algorithms and, in some cases, approaches the performance of AlphaZero.
@@
The results of this study show that AlphaViT, AlphaViD, and AlphaVDA outperform traditional methods such as Minimax and MCTS across all tested scenarios. Although AlphaZero remains the top performer in some cases, particularly for games with larger boards, the proposed agents exhibit competitive performance. In Othello, AlphaViT, AlphaViD with a deep DNN (AlphaViD L5), and AlphaVDA with a deep DNN (AlphaVDA L5) achieve Elo ratings equal to or higher than AlphaZero. Furthermore, multigame-trained variants perform on par with or better than single-game-trained variants.
@@
AlphaViT, AlphaViD, and AlphaVDA show strong adaptability across different games and board sizes. The agents with DNNs trained on a single game often achieve performance comparable to traditional game algorithms, such as Minimax and MCTS, even when playing on board sizes on which they are not trained. Moreover, multi-game-trained agents frequently perform on par with or surpass their single-game-trained counterparts. The agents with fine-tuned DNNs trained on small-board games consistently achieve performance equal to or better than that of agents with non-fine-tuned DNNs. In the case of Gomoku, using pre-trained weights from small-board games significantly accelerates convergence and enhances the final performance. This suggests effective knowledge transfer between different board sizes, mirroring human learning processes which skills acquired in simpler game variants (e.g., 9x9 Go) can be applied to more complex versions (e.g., 19x19 Go). Such adaptability suggests that the proposed agents may have significant potential for advancing the development of multitask AI.
@@
Comparing the three proposed agents reveals that AlphaViT L4 outperforms both AlphaViD L1 and AlphaVDA L1, despite having an equivalent number of parameters. This difference in performance may be attributed to the fewer transformer encoder layers in AlphaViD and AlphaVDA. This observation is further supported by the fact that AlphaViD L5 and AlphaVDA L5 exhibit performance comparable to AlphaViT L4. The simpler architecture of AlphaViT, consisting solely of encoder layers, may lead to more efficient performance in certain games, even when it has fewer parameters than its variants. However, this simplicity constrains AlphaViT’s flexibility, as its output size is fixed to the number of input embeddings, limiting its applicability to games beyond the classic board games. In contrast, including a decoder layer in AlphaViD and AlphaVDA allows for dynamic adjustment of the policy vector size, providing greater adaptability to games with varying action spaces. This architectural flexibility makes AlphaViD and AlphaVDA versatile candidates for handling more complex games or environments with continuous action spaces.
@@
Our results also relate to those of Soemers et al. (2023), who highlighted the importance of transferring trained policies and value functions across games with varying board sizes and action spaces. Similarly, our agents, equipped with DNNs that utilize weights fine-tuned from a small-board game or trained simultaneously on multiple games, demonstrate enhanced gameplay skills. This suggests that knowledge obtained from small-board games and other games is used efficiently. The ability to train on multiple games concurrently may also help these agents avoid overfitting, further enhancing their generalization capabilities.
@@
Although our proposed agents with transformer-based DNNs are slightly weaker than AlphaZero, we believe they have the potential for further improvement through enhanced training techniques. For example, an efficient decay of the learning rate during training iterations (Silver et al., 2016, 2016, 2018) and warm-start method (Wang et al., 2020, 2021) will improve performance. Increasing the number of self-plays per iteration may also be effective.

### v3

@@ 5.2 Variation of Elo rating over training iteration
Setup.
Figures 4 and 5 illustrate the progression of Elo ratings for AlphaViT, AlphaViD, AlphaVDA, and AlphaZero across multiple training iterations for large and small board configurations in three games: Connect 4, Gomoku, and Othello. Elo ratings were recorded at every iteration from 1 to 10, subsequently every 20 iterations up to iteration 100, and thereafter at intervals of 100 iterations until iteration 3000 for large board configurations and 1000 for small board configurations. The shaded areas around each Elo curve indicate the 95% confidence intervals computed through bootstrapping, as detailed in Appendix F. One iteration corresponds to a complete cycle of self-play data generation, data augmentation, and a single update of the neural network (see Appendix B). The opponents used for evaluation were identical to those described in previous experiments: AlphaViT L4 LB, AlphaViD L1 LB, AlphaVDA L1 LB, AlphaViT L4 SB, AlphaViD L1 SB, AlphaVDA L1 SB, AlphaViT L4 Multi, AlphaViD L1 Multi, AlphaVDA L1 Multi, AlphaZero, MCTS400, MCTS100, Minimax, and Random agents. The Elo ratings of these opponents were fixed as listed in Tables 3, 4, and 5. Evaluated agents’ Elo ratings were initialized to 1500 and then calculated through 40-game matches against each opponent, consisting of 20 games as the first player and 20 games as the second player.
Large board configurations.
Across all tested large board games, AlphaViT, AlphaViD, and AlphaVDA variants demonstrate rapid Elo improvement during the initial training phase, achieving approximately 80–90 % of their maximum Elo within the first 300–500 iterations. In Gomoku, the single-task agents peak near 1000 iterations and then exhibit a moderate decline. In Othello, the growth rate slows markedly after roughly 300 iterations, although the Elo ratings continue to increase slightly thereafter. Multitask-trained agents (AlphaViT Multi, AlphaViD Multi, AlphaVDA Multi) show learning speeds comparable to those of single-task models, confirming that multitask training does not adversely affect initial convergence. Notably, single-game-trained Gomoku models (AlphaViT L4 LB, AlphaViD L1 LB, AlphaVDA L1 LB) exhibit moderate Elo declines after reaching peak performance, indicating potential overfitting in later training stages. Conversely, multitask-trained Gomoku models (AlphaViT L4 Multi, AlphaViD L1 Multi, AlphaVDA L1 Multi) remain stable without a pronounced decline, highlighting multitask learning as a potentially effective approach for mitigating overfitting.
Small board configurations.
[…]
Insights and implications.
These findings clearly demonstrate that AlphaViT L4, AlphaViD L1, and AlphaVDA L1 quickly achieve high-performance levels on small boards. However, achieving similar performance stability on larger board configurations remains more challenging. The observed plateau in Elo improvement on large boards implies that simply extending training iterations beyond a certain point (around 1000 iterations) yields limited additional benefit. Therefore, to achieve further performance improvements on larger boards, enhancing model complexity—such as employing deeper transformer encoders—rather than solely prolonging training duration may be necessary.
@@
Figure 4: Elo rating progression over training iterations for AlphaViT, AlphaViD, and AlphaVDA in large board configurations (Connect 4, Gomoku, and Othello). The left column represents single-game-trained agents (single-task agents), whereas the right column indicates multi-game-trained agents (multitask agents). Solid lines represent Elo ratings calculated directly from aggregated game outcomes without employing bootstrapping, while shaded bands correspond to the 95% confidence intervals around these Elo ratings.
Figure 5: Elo rating progression over training iterations for AlphaViT, AlphaViD, and AlphaVDA in their small board configurations (Connect 4 5x4, Gomoku 6x6, and Othello 6x6). Solid lines represent Elo ratings calculated directly from aggregated game outcomes without employing bootstrapping, while shaded bands correspond to the 95% confidence intervals around these Elo ratings.
@@ 5.3 Effect of transformer encoder depth on performance
In this subsection, we evaluated the performance of the agents with deeper encoders: AlphaViT L8, AlphaViD L5, and AlphaVDA L5. Figures 6 and 7 show the variations in Elo ratings for these agents over the iterations on large and small boards, respectively. All experiments followed the evaluation protocol described in the Setup paragraph of Section 5.2.
For large board configurations, as illustrated in Figure 6, the Elo ratings of agents with deeper encoders gradually stabilize between roughly 1000 and 2000 iterations. This trend indicates a more protracted improvement phase compared to their shallower counterparts (AlphaViT L4, AlphaViD L1, and AlphaVDA L1). Conversely, in small board configurations, depicted in Figure 7, Elo ratings, as well as their 95% CIs, converge more rapidly, similar to those of the shallow encoder agents. This suggests that in less complex game environments, the additional depth of encoders offers little substantial benefit.
Table 6 lists mean Elo ratings averaged over iterations 2100–3000 for large board configurations. This highlights the performance gains achieved through the increased encoder depth, with AlphaViD and AlphaVDA demonstrating the most notable improvements. For example, AlphaViD L5 Multi and AlphaVDA L5 Multi achieve gains of +286+286+ 286 and +231+231+ 231 Elo points, respectively, in Gomoku, indicating significant performance enhancement.
The ratio of Elo ratings between the deeper and baseline DNNs, illustrated in Figure 8, further substantiates these findings. Across all evaluated games, deeper architectures generally exhibit superior performance, with AlphaViD Multi and AlphaVDA Multi achieving the highest ratios. Specifically, in Gomoku, these agents achieve Elo ratios of 1.177 and 1.149, respectively. They also outperform the baseline models in Connect 4, with Elo ratios of 1.108 and 1.130, respectively. While single-game-trained AlphaViD and AlphaVDA show marked gains in Connect 4 and Othello, their performance improvements in Gomoku are modest. In contrast, both single- and multi-game-trained AlphaViT variants exhibit relatively weaker improvements than AlphaViD and AlphaVDA. In conclusion, these results collectively demonstrate that increasing the depth of the transformer encoder layers positively influences agent performance across various game types and board configurations. This effect is particularly pronounced for AlphaViD and AlphaVDA, especially in larger and more complex game settings.
Figure 6: Elo trajectories for the deep configurations on standard boards. The left column represents single-game-trained agents, whereas the right column indicates multi-game-trained agents. Solid lines represent Elo ratings calculated directly from aggregated game outcomes without employing bootstrapping, while shaded bands correspond to the 95% CIs around these Elo ratings.
Figure 7: Elo trajectories for the deep configurations on small boards. Solid lines represent Elo ratings calculated directly from aggregated game outcomes without employing bootstrapping, while shaded bands correspond to the 95% CIs around these Elo ratings.
@@
Game Connect 4 ($\Delta$ Elo) Gomoku ($\Delta$ Elo) Othello ($\Delta$ Elo)
AlphaZero 1955(-) 2019 (-) 1925 (-)
AlphaViT L4 LB 1870 (-) 1703 (-) 1890 (-)
AlphaViT L8 LB 1915 (+45) 1804 (+101) 1942 (+52)
AlphaViD L1 LB 1747 (-) 1630 (-) 1726 (-)
AlphaViD L5 LB 1851 (+104) 1682 (+52) 1843 (+117)
AlphaVDA L1 LB 1738 (-) 1569 (-) 1749 (-)
AlphaVDA L5 LB 1894 (+156) 1596 (+27) 1884 (+135)
AlphaViT L4 Multi 1827 (-) 1835 (-) 1848 (-)
AlphaViT L8 Multi 1909 (+82) 1947 (+112) 1883 (+35)
AlphaViD L1 Multi 1658 (-) 1616 (-) 1669 (-)
AlphaViD L5 Multi 1838 (+180) 1902 (+286) 1704 (+35)
AlphaVDA L1 Multi 1666 (-) 1556 (-) 1670 (-)
AlphaVDA L5 Multi 1884 (+218) 1787 (+231) 1833 (+163)
Mean Elo ratings from 2100 to 3000 iterations for each agent. $\Delta$ Elo indicates the difference from a shallower model. Bolded ratings indicate agents whose Elo rating is within 100 points of AlphaZero.
Figure 8: The ratio of Elo ratings between deeper and baseline DNNs for AlphaViT, AlphaViD, and AlphaVDA across Connect 4, Gomoku, and Othello. Ratios greater than 1.0 indicate superior performance of agents with deeper DNNs compared to those with baseline DNNs. The error bars represent 95% CIs calculated through bootstrapping (see Appendix F).
@@ 5.4 Effect of fine-tuning from small board games
Figure 9 illustrates the Elo ratings of AlphaViT L8, AlphaViD L5, and AlphaVDA L5 with fine-tuned and non-fine-tuned (randomly initialized) DNNs across three games: Connect 4, Gomoku, and Othello. The fine-tuned DNNs were initialized using weights from the single-task DNNs that had been trained for 200 iterations on the small board configuration. For example, the fine-tuned DNN for Connect 4 was initialized with weights from the DNN trained on Connect 4 5x4. In contrast, agents with non-fine-tuned DNNs were initialized with random weights. All experiments followed the evaluation protocol described in the Setup paragraph of Section 5.2. For agents with non-fine-tuned (randomly initialized) DNNs, we reused the Elo ratings previously reported in Section 5.3.
@@
The results demonstrate that agents with fine-tuned DNNs consistently achieve higher Elo ratings than those with randomly initialized DNNs for Gomoku and Othello. The performance difference is especially pronounced during the early iterations; in this phase, fine-tuned DNNs improve rapidly and then stabilize for Gomoku and Othello. For Gomoku, the advantage of fine-tuning is clear, as the agents with fine-tuned DNNs consistently outperform those with non-fine-tuned DNNs throughout the training process. The plateau appears in AlphaViD L5 and AlphaVDA L5 for Gomoku between approximately 1000 and 2000 iterations. However, the performance improves after 2000 iterations. For Connect 4, fine-tuning yields little or no improvement in Elo.
@@
Figure 9: Elo rating progression over training iterations of the agents (AlphaViT L8, AlphaViD L5, and AlphaVDA L5) with fine-tuned and randomly initialized DNNs for three board games (Connect 4, Gomoku, and Othello). The three columns show the results for AlphaViT L8, AlphaViD L5, and AlphaVDA L5, respectively. The three rows show the results for Connect 4, Gomoku, and Othello. The shaded bands represent the 95% confidence intervals around the Elo ratings, calculated through bootstrapping (see Appendix F).
6 Conclusion
@@
In this paper, we present AlphaViT, AlphaViD, and AlphaVDA, which are novel game-playing AI agents designed to overcome the limitations of AlphaZero using Vision Transformers (ViT). Unlike AlphaZero, which is restricted to fixed board sizes, our proposed agents demonstrate adaptability, handle different board sizes effectively, and exhibit flexibility across games. Furthermore, these agents can simultaneously train on and play multiple games, such as Connect 4, Gomoku, and Othello, within a single shared neural network. The performance of these multitask agents surpasses that of traditional game AI algorithms and, in some cases, approaches AlphaZero’s performance.
@@
The results of this study show that AlphaViT, AlphaViD, and AlphaVDA outperform traditional methods such as Minimax and MCTS across all tested scenarios. Although AlphaZero remains the top performer in some cases, particularly for games with larger boards, the proposed agents exhibit competitive performance. AlphaViT L8 matches AlphaZero in Connect 4 and Othello. In Othello, the deeper versions (L5) of AlphaViD and AlphaVDA narrow the gap with AlphaZero but do not yet surpass it. Multigame-trained variants perform on par with or better than single-game-trained variants with deeper DNNs in Connect 4 and Gomoku, while remaining slightly behind in Othello.
@@
AlphaViT, AlphaViD, and AlphaVDA show strong adaptability across different games and board sizes. The agents with DNNs trained on a single game often achieve performance comparable to traditional game algorithms, such as Minimax and MCTS, even when playing on board sizes on which they are not trained. Moreover, multi-game-trained agents frequently perform on par with or surpass their single-game-trained counterparts. The agents with fine-tuned DNNs trained on small board games achieve performance better than that of agents with non-fine-tuned DNNs in Gomoku and Othello, but show little or no improvement in Connect 4. In the case of Gomoku, pre-trained weights from small board games significantly accelerate convergence and enhance the final performance. This suggests effective knowledge transfer between different board sizes, mirroring human learning processes, in which skills acquired in simpler variants (e.g., 9x9 Go) can be applied to more complex versions (e.g., 19x19 Go). Such adaptability suggests that the proposed agents may have significant potential for advancing the development of multitask AI.
@@
Comparing the three proposed agents reveals that AlphaViT L4 outperforms both AlphaViD L1 and AlphaVDA L1, despite having a similar number of parameters. This difference in performance may be attributed to the smaller number of encoder layers used in AlphaViD and AlphaVDA. This observation is further supported by the fact that AlphaViD L5 and AlphaVDA L5 exhibit performance comparable to AlphaViT L4. The simpler architecture of AlphaViT, consisting solely of encoder layers, may lead to more efficient performance in certain games, even when it has fewer parameters than its variants. However, this simplicity constrains AlphaViT’s flexibility, as its output size is fixed to the number of input embeddings, limiting its applicability to games beyond classic board games. In contrast, including a decoder layer in AlphaViD and AlphaVDA allows for dynamic adjustment of the policy vector size, providing greater adaptability to games with varying action spaces. This architectural flexibility makes AlphaViD and AlphaVDA versatile candidates for handling more complex games or environments with continuous action spaces.
7 Discussion
@@
Our results also relate to those of Soemers et al., (2021, 2023), who highlighted the importance of transferring trained policies and value functions across games with varying board sizes and action spaces. Similarly, our agents, equipped with DNNs that utilize weights either fine-tuned from a small board game or trained simultaneously on multiple games, demonstrate enhanced gameplay skills. This suggests that knowledge obtained from small board games and other games is used efficiently. The ability to train on multiple games concurrently may also help these agents avoid overfitting, further enhancing their generalization capabilities.
As shown in Tables 3, 4, and 5, AlphaZero achieves the top performance with the smallest number of parameters in some cases, particularly for games with larger boards. This likely stems from the strong inductive bias inherent in its convolutional and residual network architecture, which is well-suited for single games and enables it to easily extract local and spatial features (He et al.,, 2016; Battaglia et al.,, 2018). Remarkably, this architecture also allows some flexibility in board size when game rules remain unchanged. For example, Wu’s KataGo preserves the convolutional backbone while introducing global pooling layers—calculating channel-wise mean, scaled mean, and maximum—to both the trunk and heads, enabling a unified network to play Go from 9x9 through 19x19 without increasing the number of parameters (Wu,, 2019).
However, the strong inductive biases of CNNs may limit adaptability when game rules or mechanics significantly change. Soemers et al., (2021) systematically demonstrated that fully convolutional networks can effectively transfer learning across different board sizes or minor variants of the same game but struggle significantly when transferring between games with substantially different rules or mechanics, resulting in poor or negative transfer performance even after fine-tuning. In contrast, transformer-based agents sacrifice some of this inductive bias in favor of greater flexibility (Dosovitskiy et al.,, 2021; d’Ascoli et al.,, 2022). This enables them to handle multiple games and variable board sizes at the cost of an increase in parameter count. In this study, when we employed deep transformer encoders and model parameters of AlphaViT, AlphaViD, and AlphaVDA were much larger than those of AlphaZero’s DNN, their performance approached, and in Othello even surpassed, that of AlphaZero. This trade-off highlights a key design consideration: models with strong inductive bias can achieve higher efficiency in specialized domains, whereas more general architectures require additional capacity to compensate for their weaker assumptions (d’Ascoli et al.,, 2022).
@@
Although our proposed agents with transformer-based DNNs are slightly weaker than AlphaZero, we hypothesize they have the potential for further improvement through enhanced training techniques. For example, properly scheduled learning-rate decay during training iterations (Silver et al.,, 2016, 2017, 2018) and warm-start method (Wang et al.,, 2020, 2021) will improve performance. Increasing the number of self-plays per iteration may also be effective.
All experiments were carried out on affordable, consumer-grade GPUs (RTX 4060 Ti 16 GB and RTX 3060 12 GB) rather than on datacenter accelerators. Due to the significant memory requirements for training transformer architectures, even moderately sized boards, such as Gomoku 9x9 and Othello, approached the memory limits of these GPUs. To address this issue, we enabled automatic mixed-precision training with torch.cuda.amp.autocast(), so that most tensor operations run in FP16/BF16, while numerically sensitive layers and master weights remain in FP32. Additionally, we employed data-parallel training across multiple GPUs within a single custom-built PC to further reduce per-device memory usage. These practical choices emphasize the reproducibility of AlphaViT but also restrict experimentation with deeper transformer encoders and larger boards, such as 19x19 Go. Nevertheless, the proposed architecture inherently supports variable input sizes, and our fine-tuning experiments (Section 5.4) indicate that weights pre-trained on small board games effectively generalize to larger board configurations. Consequently, when more powerful hardware is available, it will be feasible to first train agents with larger transformer encoders on small boards and then efficiently fine-tune, rather than training from scratch, to extend the proposed agents to full-sized boards and more complex games.
@@
Acknowledgments
The author gratefully acknowledges the assistance of large language models in improving the grammar and style of this work.